Quartiles & Box plots:
Insert your student number into column A. Multiply column A by column B and place the result in column C. (my student number is 2014120)
A B C
* 1 =
* 2 =
* 3 =
* 4 =
* 5 =
* 6 =
* 7 =
Column C is now Data Set 1.
Tasks:
C.1
Produce a box plot for Data Set 1.
C.2
Calculate the Interquartile Range for Data Set 1.




Answer by jim_thompson5910(35256)


A*BC = A*B
2*12
0*20
1*33
4*416
1*55
2*612
0*70


Data set 1 is this list (from column C): 2,0,3,16,5,12,0

========================================================
C.1
Box Plot (I used geogebra to make this boxplot)



Five number summary
min = 0
Q1 = 0
median = 3
Q3 = 12
max = 16

-------------------------------------------------------
C.2

Interquartile Range (IQR)

IQR = Q3 - Q1
IQR = 12 - 0
IQR = 12

This is the distance from edge to edge of the box.